A body of mass m is thrown at an angle α to the horizontal with the initial velocity v0. Find the mean power developed by gravity over the whole time of motion of the body, and the instantaneous power of gravity as a function of time.
1
Expert's answer
2020-10-16T11:00:24-0400

Solution

The velocity of body after t sec is given by

"\overrightarrow{ v}=\overrightarrow{ v_0}+ \overrightarrow{g} t"

Power is given by

"P=\overrightarrow{F}. \overrightarrow{v}=m\overrightarrow{g}. \overrightarrow{v}"

Here (.) denotes dot product.

"P= m \overrightarrow{g} .( \overrightarrow{ v_0}+ \overrightarrow{g} t)"

"P=mg(gt-v_0 \sin \alpha)"

Then mean power

"<P>=\frac{E}{\tau}=\frac{m \overrightarrow{g} .\Delta \overrightarrow{ r}}{\tau}"

Because mg and "\Delta r" are perpendicular to each other so

Mean power

<P>=0